OpenStudy (anonymous):

solve the system of inequalities y > 2x - 1 and 3x + 2y = 10

OpenStudy (anonymous):

@jim_thompson5910 Can you please help?

jimthompson5910 (jim_thompson5910):

is there a typo on 3x + 2y = 10 ?

OpenStudy (anonymous):

yea!, 3x + 2y < 10

jimthompson5910 (jim_thompson5910):

is there a line under any of the inequalities?

OpenStudy (anonymous):

nope

jimthompson5910 (jim_thompson5910):

the first task is to graph y > 2x - 1

jimthompson5910 (jim_thompson5910):

you do this by step 1) graphing the line y = 2x - 1 step 2) make this a dashed line step 3) shade above the dashed line

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

Yea, whats next

jimthompson5910 (jim_thompson5910):

graph the second inequality

OpenStudy (anonymous):

I dont know how

OpenStudy (anonymous):

hello?

jimthompson5910 (jim_thompson5910):

graphing 3x + 2y < 10 step 1) graph 3x + 2y = 10 step 2) make this a dashed line step 3) shade below the dashed line

jimthompson5910 (jim_thompson5910):

let me know if that helps or not

OpenStudy (anonymous):

I dont know how to graph the second equation though

OpenStudy (anonymous):

can you show me?

jimthompson5910 (jim_thompson5910):

how about we solve for y

jimthompson5910 (jim_thompson5910):

3x + 2y = 10 2y = -3x + 10 y = -3x/2 + 10/2 y = -3/2x + 5

jimthompson5910 (jim_thompson5910):

does that help?

OpenStudy (anonymous):

yes, very much so, :) thanks

jimthompson5910 (jim_thompson5910):

ok great, glad it does